OpenStudy (anonymous):

Need to see if someone can help me to see if I did this correctly ty. Solve for y using the addition principle: = -1/8 + y= -3/4 -1/8 + y +1/8= -3/4 + 1/8 -1/8 + 1/8 = ( 1)/( 8) - 1/8= 0 y= -3/4 + 1/8 y= -7/8

OpenStudy (anonymous):

Umm not quite.

OpenStudy (anonymous):

OK what did I not do?

OpenStudy (anonymous):

-(1)/(8)+y=-(3)/(4) Reorder the polynomial -(1)/(8)+y alphabetically from left to right, starting with the highest order term. y-(1)/(8)=-(3)/(4) Since -(1)/(8) does not contain the variable to solve for, move it to the right-hand side of the equation by adding (1)/(8) to both sides. y=(1)/(8)-(3)/(4) To add fractions, the denominators must be equal. The denominators can be made equal by finding the least common denominator (LCD). In this case, the LCD is 8. Next, multiply each fraction by a factor of 1 that will create the LCD in each of the fractions. y=-(3)/(4)*(2)/(2)+(1)/(8) Multiply -3 by 2 to get -6. y=-(6)/(2*4)+(1)/(8) Multiply 4 by 2 to get 8. y=-(6)/(8)+(1)/(8) Combine the numerators of all fractions that have common denominators. y=(-6+1)/(8) Add 1 to -6 to get -5. y=(-5)/(8) Move the minus sign from the numerator to the front of the expression. y=-(5)/(8)

OpenStudy (anonymous):

@bonjovifan03 Do you see what you missed now? Did you understand now? Any questions/comments just ask me.

hartnn (hartnn):

the only thing that should have been shown here, was y= -3/4 + 1/8 y= -7/8 <----incorrect i would like to know from @bonjovifan03 how she got -7/8 ? it might be just a silly mistake of -6+1 = -7...

hartnn (hartnn):

as everything else was correct.

OpenStudy (anonymous):

You are both correct. I made the mistake like hartnn said. I appreciate both your input. Since I am not good with Algebra as you can tell..

hartnn (hartnn):

the more you practice, the better you become .... :)

OpenStudy (anonymous):

its okaii we are help to help you :)

OpenStudy (anonymous):

ty again...